OpenStudy (grant330sims):

Let f(x) = 3x – 42 and g(x) = 5x. Part 1 [4 points] What is (f ◦ g)(x) ? Part 2 [2 points] Use complete sentences to describe the method you used to solve this problem.

OpenStudy (lalaly):

(f ◦ g)(x) =f(g(x)) f(5x)=3(5x)-42=15x-42

OpenStudy (grant330sims):

sweet, but how exactly did you do it?

OpenStudy (lalaly):

well g(x) =5x right?

OpenStudy (grant330sims):

right.

OpenStudy (lalaly):

ok so when u say f(g(x)) that means f(5x) and f(5x) is simply substituting the expression g(x) in x in f(x) f(x)=3x-42 f(g(x))=3(g(x))-42 f(5x)=3(5x)-42

OpenStudy (grant330sims):

Thank you so much!

OpenStudy (lalaly):

no problem:D